suppose a uniform spherical star of mass M and radius R collapses to a uniform sphere of radius 10^-5 .If the original star had a rotation rate of 1 rev each 25 days(as does the sun) ,what will be the rotation rate of the resulting object
1
Expert's answer
2020-04-20T15:07:41-0400

Angular Momentum Conservation: In the absence of an external torque, the net angular momentum of a system remains conserved.

"L\n=\nI\n\u03c9\n=\nc\no\nn\ns\nt\n \n\u21d2\nI_\n1\n\u03c9_\n1\n=\nI_\n2\n\u03c9_\n2"


The Moment-of-Inertia of a sphere of mass M and radius R spinning along an axis passing through its centre is :"I
=
\frac2
5
M
R^
2"

now applying momentum conservation

"\frac2
5
M
R^
2\times\frac{2\pi}{25}=
\frac2
5
M
(10^{-5}R)^
2\times w_2"

"w_2=8\pi\times10^8"